Find the rate of change of area of the square in m/s with respect to its side when the sides, s, is 4m each.

1
Expert's answer
2021-10-15T05:45:55-0400
"A=A(s)=s^2"

"A'=(s^2)'=2s"

"A'(4)=2(4)=8"

When "s=4\ m," the rate of change of area is "8\ m^2/m."